Light travelling from water into a diamond, strikes the diamond at an angle of 42.0 degrees. What is the angle of refraction?



1
Expert's answer
2020-12-31T08:08:36-0500

The angle of refraction (see Snell's law) is:


"\sin\theta_r = \dfrac{n_1}{n_2}\sin\theta_i"

where "n_2 = 2.417" is the refractive index of diamond, "n_1 = 1.333" is the refractive index of water, and "\theta_i = 42\degree" is the incident angle. Thus, obtain:


"\theta_r = \arcsin\left(\dfrac{1.333}{2.417}\sin42\degree\right) \approx 21.7\degree"

Answer. "21.7\degree".